# Univariate-Guided Interaction Modeling

### v1

@@ 2.1 Setup and Notation
$\mathcal{P}=\{(j,k)\in[p]^{2}:\ j<k\}\quad\text{and}\quad Z=\big(X_{j}\odot X_{k}\big)_{(j,k)\in\mathcal{P}}\in\mathbb{R}^{n\times\binom{p}{2}}$
@@ 7.1 The UniLasso step
    Assume that $|S_{M}|$, $|S_{I}|$, $|\beta_{0,M}^{*}|$, $|\beta_{0,I}^{*}|$, $\max_{j\in S_{M}^{\prime}}|\beta_{j}^{*}|$, $\max_{(j,k)\in S_{I}}|\gamma_{jk}^{*}|$, $\max_{j\in S_{M}^{\prime}}\mathbb{E}(|X_{ij}|)$ and $\max_{j\in[p]}|\mathrm{Cov}(X_{ij},Y_{i})|$ are all upper bounded by a positive absolute constant $0<C_{9}<\infty$.
@@
$\mathbb{P}\Big(\forall j\notin S_{M}^{\prime}\ \widehat{\beta}_{j}^{s}=0,\max_{j\in S_{M}}|\widehat{\beta}_{j}^{s}-\beta^{*}_{j}|\leq C\lambda_{1},|\widehat{\beta}_{0}^{s}-\beta^{*}_{0,M}|\leq C\lambda_{1}\Big)$ (3)
$\geq 1-Cpn\,\exp(-cn^{3/5}\lambda_{1}^{2})-Cn\,\exp(-cn^{3/5})\xrightarrow[n\to\infty]{}1$ (4)
@@
    $B^{\prime}=\max\big\{\max_{(j,k)\notin S_{I}}|\beta_{1,jk}^{*,\mathrm{uni}}|,\max_{j\notin S_{M}}|\beta_{1,j}^{*,\mathrm{uni}}|\big\}.$
@@
    Assume that $|S_{M}|$, $|S_{I}|$, $\max_{j\in S_{M}^{\prime}}|\beta_{j}^{*}|$, $\max_{(j,k)\in S_{I}}|\gamma_{jk}^{*}|$, $\max_{j\in S_{M}^{\prime}}\mathbb{E}(|X_{ij}|)$, $\max_{(j,k)\in S_{I}}\mathbb{E}(|X_{ij}X_{ik}|)$ and $\max_{(j,k)\in\mathcal{P}}|\mathrm{Cov}(X_{ij}X_{ik},Y_{i})|$ are all upper bounded by a positive absolute constant $0<C_{16}<\infty$.
@@
$\mathbb{P}\Big(\forall j\notin S_{M}^{\prime}\ \widehat{\beta}_{j}^{s}=0,\forall(j,k)\notin S_{I}\ \widehat{\beta}_{jk}^{s}=0,\max_{j\in S_{M}^{\prime}}|\widehat{\beta}_{j}^{s}-\beta^{*}_{j}|\leq C^{\prime}\lambda,\max_{(j,k)\in S_{I}}|\widehat{\beta}_{jk}^{s}-\gamma^{*}_{jk}|\leq C^{\prime}\lambda\Big)$ (6)
$\geq 1-C^{\prime}p^{2}n\,\exp(-c^{\prime}n^{3/5}\lambda_{1}^{2})-C^{\prime}n\,\exp(-c^{\prime}n^{1/3})\xrightarrow[n\to\infty]{}1$ (7)
@@ 7.2 Triplet Scans and largest log-gap rule
Let $\widehat{\beta}_{jk,jk}$ be the estimated coefficient for $\beta_{jk,jk}$ in 9. CSIS uses $|\widehat{\beta}_{jk,jk}|$ as the screening statistic for the interaction $X_{j}\odot X_{k}$. The screening rule is defined by thresholding the magnitude of the conditional marginal coefficient, ie
@@ 8 Discussion
$\mathcal{P}=\{(j,k)\in[p]^{2}:\ j<k\}\quad\text{and}\quad Z=\big(X_{j}\odot X_{k}\big)_{(j,k)\in\mathcal{P}}\in\mathbb{R}^{n\times\binom{p}{2}}$
@@
Under A1, we have $\|X_{i}\|_{\psi_{2}}<\infty.$ By Lemma 2.7.7, Exercise 2.7.10 and Example 2.5.8 Vershynin [11], we get
@@
$|X_{ij}X_{ik}X_{ir}|^{2/3}\leq\frac{1}{3}X_{ij}^{2}+\frac{1}{3}X_{ik}^{2}+\frac{1}{3}X_{ir}^{2}\quad\text{and}\quad|X_{ij}X_{ik}X_{ir}X_{im}|^{1/2}\leq\frac{1}{4}X_{ij}^{2}+\frac{1}{4}X_{ik}^{2}+\frac{1}{4}X_{ir}^{2}+\frac{1}{4}X_{im}^{2}$
we get that $X_{ij}X_{ik}X_{ir}$ is sub-Weibull(2/3) and $X_{ij}X_{ik}X_{ir}X_{im}$ is sub-Weibull(1/2). Also, $Y_{i}^{2}$ is sub-Weibull(1/2) because $\|Y_{i}\|_{\psi_{1}}<\infty$, and $X_{ij}^{2}$ is sub-exponential because $\|X_{ij}\|_{\psi_{2}}<\infty$. $X_{i}$ is sub-gaussian so in particular is sub-Weibull(2/3). Hence,
@@
  • $X_{ij}^{2}$, $Y_{i}$ and $X_{ij}X_{ik}$ are sub-exponential, so by Bernstein’s inequality,
    $\mathbb{P}\!\left(\big|P_{n}[X_{ij}^{2}]-\mathbb{E}[X_{ij}^{2}]\big|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,n\,\min\!\Big\{\tfrac{t^{2}}{\|X_{ij}^{2}\|_{\psi_{1}}^{2}},\tfrac{t}{\|X_{ij}^{2}\|_{\psi_{1}}}\Big\}\right)$
@@
  • $X_{ij}$ is sub-Gaussian, so by Hoeffding’s inequality,
@@
From A1, A2, and A7 and using our previous results that control the $\|.\|_{\psi}$ of the different quantities of interest, we see in particular that $\|\Gamma^{*}\|_{L^{1}},\|\beta^{*}\|_{2}$ are $O_{a}(1)$ and so are $\|X_{ij}X_{ik}\|_{\psi_{1}}$, $\|Y_{i}\|_{\psi_{1}}$, $\|X_{ij}^{2}\|_{\psi_{1}}$, $\|\epsilon_{i}^{\prime}X_{il}\|_{\psi_{2/3}}$, $\|X_{ij}Y_{i}\|_{\psi_{2/3}}$, $\|X_{ij}\|_{\psi_{2}}$, $\|X_{ij}^{2}X_{ik}^{2}\|_{\psi_{1/2}}$, and $\|Y_{i}^{2}\|_{\psi_{1/2}}$. Therefore, the concentration inequalities stated in Lemma 1 follow. $\square$
The next two lemmas are about the restricted eigenvalue concentrations.
@@
Then there exists positive absolute constants $0<K_{1},K_{2}<\infty$ such that for any $\epsilon\in(0,1)$, and all $n\geq 1$, we have :
@@
For any vector $V\in\mathbb{R}^{p}$, define $V_{S_{M}^{\prime}}=(V_{i})_{i\in S_{M}^{\prime}}\in\mathbb{R}^{|S_{M}^{\prime}|}$. By Theorem 4.6.1 and Exercise 4.7.3 in Vershynin [11], we get that for all $u\geq 0$ with probability at least $1-2\exp(-u)$
$\|P_{n}[(X_{i}-\mathbb{E}(X_{i}))_{S_{M}^{\prime}}(X_{i}-\mathbb{E}(X_{i}))_{S_{M}^{\prime}}^{T}]-\Sigma_{S_{M}^{\prime}}\|_{op}$ $\leq O_{a}(1)\frac{\|X_{i,S_{M}^{\prime}}\|_{\Psi_{2}}^{2}}{\eta_{M}^{*2}}\left(\sqrt{\frac{|S_{M}^{\prime}|+u}{n}}+\frac{|S_{M}^{\prime}|+u}{n}\right)\|\Sigma_{S_{M}^{\prime}}\|_{op}$
But $\|X_{i,S_{M}^{\prime}}\|_{\Psi_{2}}\leq\|X_{i}\|_{\Psi_{2}}$, so with probability at least $1-2\exp(-u)$
$\|P_{n}[X_{i,S_{M}^{\prime}}X_{i,S_{M}^{\prime}}^{T}]-\Sigma_{S_{M}^{\prime}}\|_{op}$ $\leq O_{a}(1)\frac{\|X_{i}\|_{\Psi_{2}}^{2}}{\eta^{*2}}\left(\sqrt{\frac{|S_{M}^{\prime}|+u}{n}}+\frac{|S_{M}^{\prime}|+u}{n}\right)\|\Sigma_{S_{M}^{\prime}}\|_{op}$
$+2\|\mathbb{E}(X_{i})_{S_{M}^{\prime}}\|_{2}\|P_{n}[X_{i,S_{M}^{\prime}}]-\mathbb{E}(X_{i})_{S_{M}^{\prime}}\|_{2}$
But for all $t>0$ we have
$\mathbb{P}\!\left(|P_{n}[X_{ij}]-\mathbb{E}[X_{ij}]|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,nt^{2}\right)$
@@
Therefore by A1, A2 and A7, we get that for all $u,t>0$ with probability at least $1-2\exp(-u)-2|S_{M}^{\prime}|\exp(-\Omega_{a}(1)nt^{2})$
$|\widehat{\eta}_{M}-\eta^{*}_{M}|\leq O_{a}(1)\left(\sqrt{\frac{|S_{M}^{\prime}|+u}{n}}+\frac{|S_{M}^{\prime}|+u}{n}+t\sqrt{|S_{M}^{\prime}|}\right)$
So for all $u>0$ such that $\frac{|S_{M}^{\prime}|+u}{n}<1$ with probability at least $1-O_{a}(1)\exp(-u)$
$|\widehat{\eta}_{M}-\eta^{*}_{M}|\leq O_{a}(1)\left(\sqrt{\frac{|S_{M}^{\prime}|+u}{n}}+\sqrt{\frac{u|S_{M}^{\prime}|}{n}}\right)$
@@
Therefore from the corresponding concentration inequalities, we get that for all $t>0$
$\mathbb{P}(|\widehat{\eta}_{A}-\eta_{A}^{*}|>|S_{A}^{\prime}|t)$ $\leq 2|S_{M}|\exp(-\Omega_{a}(1)nt^{2})+2(|S_{M}|^{2}+|S_{I}|)\exp(-\Omega_{a}(1)n\min(t,t^{2}))$ (10)
$+12|S_{M}||S_{I}|\exp(-\Omega_{a}(1)n^{\frac{3}{4}}t)+48|S_{I}|^{2}\exp(-\Omega_{a}(1)n^{\frac{3}{5}}t^{\frac{4}{5}})$ (11)
By taking $t=n^{-\epsilon/2}$ for $\epsilon\in(0,\frac{2}{3})$ and noting that $|S_{M}|,|S_{I}|=O_{a}(1)$ from A7, we get the concentration bound stated in Lemma 3. $\square$
@@
Hence, using the concentration of $L_{n,i}[X_{\ell j}^{2}]$ and $L_{n,i}[X_{\ell j}]$, we get
$\mathbb{P}(|B_{n}-b|>\frac{b}{2})\leq O_{a}(1)\exp(-\Omega_{a}(1)\,n)$
On $|B_{n}-b|\leq\frac{b}{2}$, we have $B_{n}\geq b/2=\Omega_{a}(1)$. Decompose
@@
Using the concentration inequalities for each term, we see that for $t\in(0,O_{a}(1))$
@@
Since $t\in(0,O_{a}(1))$, the second term is dominated by the first, yielding
@@
By the same concentration bounds, we get that for all $t\in(0,O_{a}(1))$
@@
The proof uses exactly the same algebraic arguments as in Proof. with the concentration rates changed according to the bounds established in 1. $\square$
@@
Fix $j\in[p]$ and assume $\widehat{\beta}_{j}^{s}>0$. Since $\widehat{\beta}_{j}^{s}=\widehat{\theta}_{j}\widehat{\beta}_{1,j}^{\mathrm{uni}}$, we get $\widehat{\theta}_{j}>0$. So $\widetilde{\nu}_{j}=0$ and we have
@@
Since $P_{n}\!\big[Y_{i}-\widehat{\theta}_{0}-\sum_{k}\widehat{\theta}_{k}\widehat{\eta}_{k}^{(-i)}\big]=0$, we get
@@
$\Big|\frac{\lambda_{1}}{2}\Big|\leq\Big(P_{n}[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j})^{2}]\Big)^{1/2}\Big(P_{n}\!\big[(Y_{i}-\widehat{\theta}_{0}-\sum_{k}\widehat{\theta}_{k}\widehat{\eta}_{k}^{(-i)})^{2}\big]\Big)^{1/2}$
@@
$P_{n}[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j})^{2}]$ $\leq 2P_{n}[X_{ij}^{2}]\,\max_{i\in[n]}(\widehat{\beta}_{1,j}^{(-i)})^{2}+4\max_{i\in[n]}(\widehat{\beta}_{0,j}^{(-i)}-\beta_{0,j}^{\star})^{2}+4(\beta_{0,j}^{\star}-\widehat{\beta}_{0,j})^{2}$
$=2P_{n}[X_{ij}^{2}]M_{j,1}^{2}+4D_{j,0}^{2}$
@@
$\mathbb{P}(\widehat{\theta}_{j}>0)$ $\leq\mathbb{P}\!\left(\frac{\lambda_{1}^{2}}{4}\leq P_{n}[Y_{i}^{2}]\,\big(2P_{n}[X_{ij}^{2}]M_{j,1}^{2}+4D_{j,0}^{2}\big)\right)$
$\leq\mathbb{P}\!\left(P_{n}[Y_{i}^{2}]\,\geq 2\mathbb{E}[Y_{i}^{2}]\right)+\mathbb{P}\!\left(P_{n}[X_{ij}^{2}]\,\geq 2\mathbb{E}[X_{ij}^{2}]\right)+\mathbb{P}\!\left(\frac{\lambda_{1}^{2}}{4}\leq 2\mathbb{E}[Y_{i}^{2}]\,\big(4\mathbb{E}[X_{ij}^{2}]M_{j,1}^{2}+4D_{j,0}^{2}\big)\right)$
@@
$+\mathbb{P}\!\left(D_{j,0}^{2}\geq\frac{\lambda_{1}^{2}}{64\mathbb{E}[Y_{i}^{2}]}\right)$
The first two terms are upper bounded respectively by $O_{a}(1)\,\exp(-\Omega_{a}(1)n^{3/5})$ and $O_{a}(1)\,\exp(-\Omega_{a}(1)n)$. By A5, $\Omega_{a}(1)|\beta_{1,j}^{*}|\leq\lambda_{1}\leq O_{a}(1)$, so
$\mathbb{P}\!\left(M_{j,1}^{2}\geq\frac{\lambda_{1}^{2}}{64\mathbb{E}[Y_{i}^{2}]\mathbb{E}[X_{ij}^{2}]}\right)$ $\leq n\mathbb{P}\!\left(|\widehat{\beta}_{1,j}^{(-i)}|\geq\Omega_{a}(1)\lambda_{1}\right)$
$\leq n\mathbb{P}\!\left(|\widehat{\beta}_{1,j}^{(-i)}-\beta_{1,j}^{*,\mathrm{uni}}|\geq\Omega_{a}(1)\lambda_{1}\right)$
@@
$\mathbb{P}\!\left(D_{j,0}^{2}\geq\frac{\lambda_{1}^{2}}{64\mathbb{E}[Y_{i}^{2}]}\right)$ $\leq n\mathbb{P}\!\left(|\widehat{\beta}_{0,j}^{(-i)}-\beta_{0,j}^{*,\mathrm{uni}}|\geq\Omega_{a}(1)\lambda_{1}\right)+\mathbb{P}\!\left(|\widehat{\beta}_{0,j}-\beta_{0,j}^{*,\mathrm{uni}}|\geq\Omega_{a}(1)\lambda_{1}\right)$
@@
The proof uses exactly the same algebraic arguments as in Proof. with the concentration rates changed according to the bounds established in 1. $\square$
@@
$\quad+\frac{\beta_{j}^{*}-\widehat{\beta}_{j}^{s}}{\beta_{1,j}^{*,\mathrm{uni}}}P_{n}\!\Big[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})\Big(\sum_{k\in S_{M}}\widehat{\theta}_{k}\big(\widehat{\beta}_{0,k}-\widehat{\beta}_{0,k}^{(-i)}+(\widehat{\beta}_{1,k}-\widehat{\beta}_{1,k}^{(-i)})X_{ik}\big)\Big)\Big]$

### v2

@@ 2.1 Setup and Notation
$\mathcal{P}=\{(j,k)\in[p]^{2}\ \,|\,\ j<k\}\quad\text{and}\quad Z=\big(X_{j}\odot X_{k}\big)_{(j,k)\in\mathcal{P}}\in\mathbb{R}^{n\times\binom{p}{2}}$
@@ 7.1 The UniLasso step
    Assume that $|S_{M}|$, $|S_{I}|$, $|\beta_{0,M}^{*}|$, $|\beta_{0,I}^{*}|$, $\max_{j\in S_{M}^{\prime}}|\beta_{j}^{*}|$, $\max_{(j,k)\in S_{I}}|\gamma_{jk}^{*}|$, $\mathbb{E}(\epsilon_{i}^{2})$ and $\max_{j\in[p]}\mathbb{E}(X_{ij}^{4})$ are all upper bounded by a positive absolute constant $0<C_{9}<\infty$.
@@
$\mathbb{P}\Big(\forall j\notin S_{M}^{\prime}\ \widehat{\beta}_{j}^{s}=0,\max_{j\in S_{M}}|\widehat{\beta}_{j}^{s}-\beta^{*}_{j}|\leq C\lambda_{1},|\widehat{\beta}_{0}^{s}-\beta^{*}_{0,M}|\leq C\lambda_{1}\Big)$
$\geq 1-Cpn\,\exp(-cn^{3/5}\lambda_{1}^{2})-Cn\,\exp(-cn^{3/5})\xrightarrow[n\to\infty]{}1$
@@
    $B^{\prime}=\max\big\{\max_{(j,k)\notin S_{I}}|\beta_{1,jk}^{*,\mathrm{uni}}|,\max_{j\notin S_{M}}|\beta_{1,j}^{*,\mathrm{uni}}|\big\}$
@@
    Assume that $|S_{M}|$, $|S_{I}|$, $\max_{j\in S_{M}^{\prime}}|\beta_{j}^{*}|$, $\max_{(j,k)\in S_{I}}|\gamma_{jk}^{*}|$, $\mathbb{E}(\epsilon_{i}^{2})$ and $\max_{j\in[p]}\mathbb{E}(X_{ij}^{4})$ are all upper bounded by a positive absolute constant $0<C_{16}<\infty$.
@@
$\mathbb{P}\Big(\forall j\notin S_{M}^{\prime}\ \widehat{\beta}_{j}^{s}=0,\forall(j,k)\notin S_{I}\ \widehat{\beta}_{jk}^{s}=0,\max_{j\in S_{M}^{\prime}}|\widehat{\beta}_{j}^{s}-\beta^{*}_{j}|\leq C^{\prime}\lambda,\max_{(j,k)\in S_{I}}|\widehat{\beta}_{jk}^{s}-\gamma^{*}_{jk}|\leq C^{\prime}\lambda\Big)$
$\geq 1-C^{\prime}p^{2}n\,\exp(-c^{\prime}n^{3/5}\lambda_{1}^{2})-C^{\prime}n\,\exp(-c^{\prime}n^{1/3})\xrightarrow[n\to\infty]{}1$
@@ 7.2 Triplet Scans and largest log-gap rule
Let $\widehat{\beta}_{jk,jk}$ be the estimated coefficient for $\beta_{jk,jk}$ in 3. CSIS uses $|\widehat{\beta}_{jk,jk}|$ as the screening statistic for the interaction $X_{j}\odot X_{k}$. The screening rule is defined by thresholding the magnitude of the conditional marginal coefficient, ie
@@ 8 Discussion
$\mathcal{P}=\{(j,k)\in[p]^{2}\ |\ \ j<k\}\quad\text{and}\quad Z=\big(X_{j}\odot X_{k}\big)_{(j,k)\in\mathcal{P}}\in\mathbb{R}^{n\times\binom{p}{2}}$
@@
Under A1, we have $\|X_{i}\|_{\psi_{2}}<\infty.$ By Lemma 2.7.7, Exercise 2.7.10 and Example 2.5.8 in Vershynin [11], we get
@@
$|X_{ij}X_{ik}X_{ir}|^{2/3}\leq\frac{1}{3}X_{ij}^{2}+\frac{1}{3}X_{ik}^{2}+\frac{1}{3}X_{ir}^{2},\qquad|X_{ij}X_{ik}X_{ir}X_{im}|^{1/2}\leq\frac{1}{4}X_{ij}^{2}+\frac{1}{4}X_{ik}^{2}+\frac{1}{4}X_{ir}^{2}+\frac{1}{4}X_{im}^{2}$
and
$|X_{ij}X_{ik}Y_{i}|^{1/2}\leq\frac{1}{4}X_{ij}^{2}+\frac{1}{4}X_{ik}^{2}+\frac{1}{2}|Y_{i}|$
we get that $X_{ij}X_{ik}X_{ir}$ is sub-Weibull(2/3), $X_{ij}X_{ik}X_{ir}X_{im}$ and $X_{ij}X_{ik}Y_{i}$ are sub-Weibull(1/2). Also, $Y_{i}^{2}$ is sub-Weibull(1/2) because $\|Y_{i}\|_{\psi_{1}}<\infty$, and $X_{ij}^{2}$ is sub-exponential because $\|X_{ij}\|_{\psi_{2}}<\infty$. $X_{i}$ is sub-gaussian so in particular is sub-Weibull(2/3) $\big(|X_{i}|^{2/3}\leq 1+|X_{i}|^{2}\big)$. Hence,
@@
  • $Y_{i}$ and $X_{ij}X_{ik}$ are sub-exponential, so by Bernstein’s inequality (Theorem 2.8.2 in Vershynin [11]),
@@
  • $X_{ij}$ is sub-Gaussian, so by Hoeffding’s inequality (Theorem 2.6.3 in Vershynin [11]),
@@
    $\mathbb{P}\!\left(\big|P_{n}[X_{ij}X_{ik}Y_{i}]-\mathbb{E}[X_{ij}X_{ik}Y_{i}]\big|>t\right)\leq 2\exp\!\left(-\Omega_{a}(1)\,n^{3/5}\,t^{4/5}/\|X_{ij}X_{ik}Y_{i}\|_{\psi_{1/2}}^{4/5}\right)$
@@
From A1, A2, and A7 and using our previous results that control the $\|.\|_{\psi}$ of the different quantities of interest, we see in particular that $\|\Gamma^{*}\|_{L^{1}},\|\beta^{*}\|_{2}$ are $O_{a}(1)$ and so are $\|X_{ij}X_{ik}X_{ir}\|_{\psi_{2/3}}$, $\|X_{ij}Y_{i}\|_{\psi_{2/3}}$, $\|\epsilon_{i}^{\prime}X_{il}\|_{\psi_{2/3}}$, $\|Y_{i}\|_{\psi_{1}}$, $\|X_{ij}X_{ik}\|_{\psi_{1}}$, $\|X_{ij}\|_{\psi_{2}}$, $\|X_{ij}X_{ik}X_{ir}X_{im}\|_{\psi_{1/2}}$, $\|X_{ij}X_{ik}Y_{i}\|_{\psi_{1/2}}$, and $\|Y_{i}^{2}\|_{\psi_{1/2}}$. Therefore, the concentration inequalities stated in Lemma 1 follow. $\square$
The next two lemmas are about the restricted eigenvalue concentrations. Denote by $\lambda_{min}(A)$ (resp. $\lambda_{max}(A)$) the smallest (resp. largest) eigenvalue of the symmetric matrix $A$.
@@
Then there exists positive absolute constants $0<K_{1},K_{2}<\infty$ such that for any $\epsilon\in(0,1)$, and all $n$ sufficiently large, we have :
@@
For any vector $V\in\mathbb{R}^{p}$, define $V_{S_{M}^{\prime}}=(V_{i})_{i\in S_{M}^{\prime}}\in\mathbb{R}^{|S_{M}^{\prime}|}$. By Theorem 4.6.1 and Exercise 4.7.3 in Vershynin [11], we get that for all $u\geq 0$ with probability at least $1-2\exp(-u)$, we have :
$\|P_{n}[X_{i,S_{M}^{\prime}}X_{i,S_{M}^{\prime}}^{T}]-\Sigma_{S_{M}^{\prime}}\|_{op}$ $\leq O_{a}(1)\max\big(1,\frac{\|X_{i,S_{M}^{\prime}}\|_{\Psi_{2}}^{2}}{\eta_{M}^{*2}}\big)\left(\sqrt{\frac{|S_{M}^{\prime}|+u}{n}}+\frac{|S_{M}^{\prime}|+u}{n}\right)\|\Sigma_{S_{M}^{\prime}}\|_{op}$
@@
Therefore, since $\|X_{i,S_{M}^{\prime}}\|_{\Psi_{2}}\leq\|X_{i}\|_{\Psi_{2}}=O_{a}(1)$ and by A1 and A3, we get that for all $u\geq 0$ with probability at least $1-2\exp(-u)$
$|\widehat{\eta}_{M}-\eta^{*}_{M}|\leq O_{a}(1)\left(\sqrt{\frac{|S_{M}^{\prime}|+u}{n}}+\frac{|S_{M}^{\prime}|+u}{n}\right)$
So for all $u>0$ such that $\frac{|S_{M}^{\prime}|+u}{n}<1$ with probability at least $1-2\exp(-u)$
$|\widehat{\eta}_{M}-\eta^{*}_{M}|\leq O_{a}(1)\sqrt{\frac{|S_{M}^{\prime}|+u}{n}}$
@@
Therefore from the corresponding concentration inequalities in 1, we get that for all $t>0$
$\mathbb{P}(|\widehat{\eta}_{A}-\eta_{A}^{*}|>|S_{A}^{\prime}|t)$ $\leq 2|S_{M}|\exp(-\Omega_{a}(1)nt^{2})+2(|S_{M}|^{2}+|S_{I}|)\exp(-\Omega_{a}(1)n\min(t,t^{2}))$
$+12|S_{M}||S_{I}|\exp(-\Omega_{a}(1)n^{\frac{3}{4}}t)+48|S_{I}|^{2}\exp(-\Omega_{a}(1)n^{\frac{3}{5}}t^{\frac{4}{5}})$
By taking $t=n^{-\epsilon/2}$ for $\epsilon\in(0,\frac{2}{3})$ and noting that $|S_{M}|,|S_{I}|,|S_{A}^{\prime}|=O_{a}(1)$ from A13, we get the concentration bound stated in Lemma 3. $\square$
@@
Hence, using the concentration of $L_{n,i}[X_{\ell j}^{2}]$ and $L_{n,i}[X_{\ell j}]$ from Lemma 1, and since $b=\Omega_{a}(1)$ from A4, we get
$\mathbb{P}(|B_{n}-b|>b/2)\leq O_{a}(1)\exp(-\Omega_{a}(1)\,n)$
On $|B_{n}-b|\leq b/2$, we have $B_{n}\geq b/2=\Omega_{a}(1)$. Decompose
@@
Using the concentration inequalities for each term from Lemma 1, we see that for $t\in(0,1)$
@@
Since $t\in(0,1)$, the second term is dominated by the first, yielding
@@
By the same concentration bounds, we get that for all $t\in(0,1)$
@@
The proof uses exactly the same algebraic arguments as in Proof. with the concentration rates changed according to the bounds established in Lemma 1. $\square$
@@
Fix $j\in[p]$ and assume $\widehat{\beta}_{j}^{s}\neq 0$. Since $\widehat{\beta}_{j}^{s}=\widehat{\theta}_{j}\widehat{\beta}_{1,j}^{\mathrm{uni}}$ and $\widehat{\theta}_{j}\geq 0$, we get $\widehat{\theta}_{j}>0$. So $\widetilde{\nu}_{j}=0$ and we have
@@
Since $P_{n}\!\big[Y_{i}-\widehat{\theta}_{0}-\sum_{k=1}^{p}\widehat{\theta}_{k}\widehat{\eta}_{k}^{(-i)}\big]=0$, we get
@@
$\Big|\frac{\lambda_{1}}{2}\Big|\leq\Big(P_{n}[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})^{2}]\Big)^{1/2}\Big(P_{n}\!\big[(Y_{i}-\widehat{\theta}_{0}-\sum_{k=1}^{p}\widehat{\theta}_{k}\widehat{\eta}_{k}^{(-i)})^{2}\big]\Big)^{1/2}$
@@
$P_{n}[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})^{2}]$ $\leq 2P_{n}[X_{ij}^{2}]\,\max_{i\in[n]}(\widehat{\beta}_{1,j}^{(-i),\mathrm{uni}})^{2}+4\max_{i\in[n]}(\widehat{\beta}_{0,j}^{(-i),\mathrm{uni}}-\beta_{0,j}^{*,\mathrm{uni}})^{2}+4(\beta_{0,j}^{*,\mathrm{uni}}-\widehat{\beta}_{0,j}^{\mathrm{uni}})^{2}$
$\leq 2P_{n}[X_{ij}^{2}]M_{j,1}^{2}+8D_{j,0}^{2}$
@@
$\mathbb{P}(\widehat{\theta}_{j}>0)$ $\leq\mathbb{P}\!\left(\frac{\lambda_{1}^{2}}{4}\leq P_{n}[Y_{i}^{2}]\,\big(2P_{n}[X_{ij}^{2}]M_{j,1}^{2}+8D_{j,0}^{2}\big)\right)$
$\leq\mathbb{P}\!\left(P_{n}[Y_{i}^{2}]\,\geq 2\mathbb{E}[Y_{i}^{2}]\right)+\mathbb{P}\!\left(P_{n}[X_{ij}^{2}]\,\geq 2\mathbb{E}[X_{ij}^{2}]\right)+\mathbb{P}\!\left(\frac{\lambda_{1}^{2}}{4}\leq 2\mathbb{E}[Y_{i}^{2}]\,\big(4\mathbb{E}[X_{ij}^{2}]M_{j,1}^{2}+8D_{j,0}^{2}\big)\right)$
@@
$+\mathbb{P}\!\left(D_{j,0}^{2}\geq\frac{\lambda_{1}^{2}}{128\mathbb{E}[Y_{i}^{2}]}\right)$
The first two terms are upper bounded respectively by $O_{a}(1)\,\exp(-\Omega_{a}(1)n^{3/5})$ and $O_{a}(1)\,\exp(-\Omega_{a}(1)n)$. By A5, if $j\notin S_{M}$, then $\Omega_{a}(1)|\beta_{1,j}^{*,\mathrm{uni}}|\leq\lambda_{1}\leq O_{a}(1)$, so by Lemma  4
$\mathbb{P}\!\left(M_{j,1}^{2}\geq\frac{\lambda_{1}^{2}}{64\mathbb{E}[Y_{i}^{2}]\mathbb{E}[X_{ij}^{2}]}\right)$ $\leq n\mathbb{P}\!\left(|\widehat{\beta}_{1,j}^{(-i),\mathrm{uni}}|\geq\Omega_{a}(1)\lambda_{1}\right)$
$\leq n\mathbb{P}\!\left(|\widehat{\beta}_{1,j}^{(-i),\mathrm{uni}}-\beta_{1,j}^{*,\mathrm{uni}}|\geq\Omega_{a}(1)\lambda_{1}\right)$
@@
$\mathbb{P}\!\left(D_{j,0}^{2}\geq\frac{\lambda_{1}^{2}}{128\mathbb{E}[Y_{i}^{2}]}\right)$ $\leq n\mathbb{P}\!\left(|\widehat{\beta}_{0,j}^{(-i),\mathrm{uni}}-\beta_{0,j}^{*,\mathrm{uni}}|\geq\Omega_{a}(1)\lambda_{1}\right)+\mathbb{P}\!\left(|\widehat{\beta}_{0,j}^{\mathrm{uni}}-\beta_{0,j}^{*,\mathrm{uni}}|\geq\Omega_{a}(1)\lambda_{1}\right)$
@@
The proof uses exactly the same algebraic arguments as in Proof. with the concentration rates changed according to the bounds established in Lemma 5. $\square$
@@
$\quad+\frac{\beta_{j}^{*}-\widehat{\beta}_{j}^{s}}{\beta_{1,j}^{*,\mathrm{uni}}}P_{n}\!\Big[(\widehat{\eta}_{j}^{(-i)}-\widehat{\beta}_{0,j}^{\mathrm{uni}})\Big(\sum_{k\in S_{M}}\widehat{\theta}_{k}\big(\widehat{\beta}_{0,k}^{\mathrm{uni}}-\widehat{\beta}_{0,k}^{(-i)}+(\widehat{\beta}_{1,k}^{\mathrm{uni}}-\widehat{\beta}_{1,k}^{(-i)})X_{ik}\big)\Big)\Big]$